# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ 7.2 Encoding equivalent classes
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq\frac{1}{2}\sum_{L=0}^{M}(2\kappa)^{M-L+R}\sum_{\ast\ast}\prod_{\operatorname{cc}\in\chi}\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}_{\operatorname{cc}}\times\mathcal{X}_{\operatorname{cc}}\end{subarray}}\prod_{f\in\{e^{\prime}\}\sqcup\mathcal{T}_{\operatorname{cc}}}(Q^{\ell(f)})_{y_{u}x_{v}}$
$\leq\frac{1}{2}\sum_{L=0}^{M}(2\kappa)^{M-L+R}\sum_{\ast\ast}N^{s-p}\delta^{L_{\mathcal{F}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell(e(\operatorname{cc}))}{Q^{\ast}}^{\ell(e^{\prime}(\operatorname{cc}))}\right)$
where we used the fact that all $\left(\mathcal{G}_{\operatorname{cc}}\right)_{\operatorname{cc}\in\chi}$ are disjoint graphs in the first line and we applied 7.7 to each $\operatorname{cc}\in\chi$ in the second line, setting $L_{\mathcal{F}}:=\sum_{\operatorname{cc}}L_{\mathcal{T}_{\operatorname{cc}}}$. Applying the same reasoning replacing $(Q^{\ell(f)})_{f}$ with ($(\{rQ\}^{\ell(f)})_{f}$ we obtain
$\sum_{\gamma^{\prime}\sim\gamma}q_{r}(\gamma^{\prime})$ $\leq\frac{1}{2}\sum_{L=0}^{M}(2r\kappa)^{M-L+R}\sum_{\ast\ast}N^{s-p}(r\delta)^{L_{\mathcal{F}}}(1-r)^{T}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell(e(\operatorname{cc}))}\{rQ^{\ast}\}^{\ell(e^{\prime}(\operatorname{cc}))}\right).$
We now sum over all $(\ell_{cc},\ell_{\operatorname{cc}}^{\prime})_{\operatorname{cc}\in\chi_{d}}$ possible and obtain
@@
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq\frac{1}{2}N^{s-p}\sum_{L=0}^{M}(2\kappa)^{M-L+R}\sum_{L_{\mathrm{d}}=0}^{L}\delta^{L-L_{\mathrm{d}}}|\mathcal{L}^{\prime}(\mathcal{F},L-L_{\mathrm{d}})|\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}},\ell_{\operatorname{cc}}^{\prime})_{\operatorname{cc}\in\chi_{\mathrm{d}}}\\ \sum_{\operatorname{cc}}\ell_{\operatorname{cc}}+\ell_{cc}^{\prime}=L_{\mathrm{d}}\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
@@
$\leq\frac{1}{2}N^{s-p}(2m(\ell+1))^{F}\sum_{L=0}^{M}\sum_{L_{\mathrm{d}}=0}^{L}(2\kappa)^{M-L+R}\delta^{L-L_{\mathrm{d}}}\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}},\ell_{\operatorname{cc}}^{\prime})_{\operatorname{cc}\in\chi_{\mathrm{d}}}\\ \sum_{\operatorname{cc}}\ell_{\operatorname{cc}}+\ell_{cc}^{\prime}=L_{\mathrm{d}}\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right).$
For all $0\leq L_{\mathrm{d}}^{1}+L_{\mathrm{d}}^{2}=L_{\mathrm{d}}$, denoting $\mathrm{x}_{\mathrm{d}}=|\chi_{\mathrm{d}}|$ we have
@@
$\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{1})\\ (\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{2})\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right)$ $\leq\sum_{\begin{subarray}{c}(\ell_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}})\\ (\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}})\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell_{\operatorname{cc}}}{Q^{\ast}}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
@@
$\leq\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{\begin{subarray}{c}(\ell_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}})\\ (\ell^{\prime}_{t})_{\operatorname{cc}}\in\mathcal{L}(T,L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}})\end{subarray}}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{\ell_{t}}{Q^{\ast}}^{\ell_{t}^{\prime}}\right)$
$\leq\Phi(a^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}{a^{\ast}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}})\leq\mathrm{N}_{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}\mathrm{N}_{L_{\mathrm{d}+\mathrm{x}_{\mathrm{d}}}^{2}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}$
where for the last two inequalities we applied 4.5 and we have used that for $b,h\in\mathcal{A}$
$\Phi(bh^{*})\leq\sqrt{\tau(bb^{*})\tau(hh^{*})},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \tau(b^{\ast}b)=\mathbb{E}[\|b^{\ast}b(e_{0})\|_{2}^{2}]\leq\|b\|^{2}.$
The integer $n$ being fixed, for all integer $L=\mathrm{q}n+\mathrm{r}\geq 1$ with $10\leq\mathrm{r}\leq n-$ 1, We set $c(n):=(\kappa\vee 1)(1+K)^{n}$ and we have:
$\mathrm{N}_{L}^{L}=\|(u+q)^{L}\|\leq\|(u+q)^{n}\|^{\mathrm{q}}\|(u+q)\|^{\mathrm{r}}\leq\mathrm{N}_{n}^{L-\mathrm{r}}(1+\|Q\|)^{n}\leq c(n)\mathrm{N}_{n}^{L}$
where we used that the operator norm is sub-multiplicative. For the convex combination case we have
@@
$(1-r)^{2(\mathrm{x}_{\mathrm{d}}-1)}\sum_{(\ell_{\operatorname{cc}},\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{\operatorname{cc}}}\{rQ^{\ast}\}^{\ell_{\operatorname{cc}}^{\prime}}\right)\leq\sum_{\ast\ast\ast\ast}(1-r)^{2(\mathrm{x}_{\mathrm{d}}-1)}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{\operatorname{cc}}}\{rQ^{\ast}\}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leq\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{\begin{subarray}{c}(\ell_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}})\\ (\ell^{\prime}_{t})_{\operatorname{cc}}\in\mathcal{L}(T,L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}})\end{subarray}}(1-r)^{2(T-1)}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{t}}\{rQ^{\ast}\}^{\ell_{t}^{\prime}}\right)$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leq\Phi(p^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}{p^{\ast}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}})\leq\mathrm{N}_{r,L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}\mathrm{N}_{r,L_{\mathrm{d}+\mathrm{x}_{\mathrm{d}}}^{2}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}.$
We obtain
@@
$\sum_{\gamma^{\prime}\sim\gamma}q(\gamma^{\prime})$ $\leq N^{s-p}(2m(\ell+1))^{F}2m(\ell+1)c(n)^{2}\sum_{L=0}^{M}\sum_{L_{\mathrm{d}}=0}^{L}(2\kappa)^{M-L+R}\delta^{L-L_{\mathrm{d}}}\mathrm{N}_{n}^{L_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}}$
$\leq N^{s-p}(2m(\ell+1))^{F}(2m(\ell+1))^{2}c(n)^{2}\rho_{n}^{2m(\ell+1)}$
@@
where we used the fact that owing to 4.7 we have $(2\kappa)\vee\delta\vee\mathrm{N}_{n}\geq 1$ and the fact that $M+R+\mathrm{x}_{\mathrm{d}}\leq 2m(\ell+1)$. It remains to apply 7.8 and adapt the constant $c(n)>0$. ∎
We now give the proof of 7.8.
Proof of 7.8.
@@
We introduce $\chi_{\mathrm{d}}^{\prime}$ the set of $\operatorname{cc}\in\chi_{\mathrm{d}}$ such that the connected component verifies $|\mathcal{X}_{\operatorname{cc}}|=1$ and there is only one incoming edge $e\in\mathrm{A}_{\gamma}$ of multiplicity $m(e)\geq 2$. For all $\operatorname{cc}\in\chi_{\mathrm{d}}^{\prime}$, we have $\mathcal{T}_{\operatorname{cc}}\backslash\{e(\operatorname{cc})\}=\emptyset$, therefore:

### v4

@@ 8.2 Encoding equivalent classes
$\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{s}}\leq\sum_{\gamma^{\prime}\sim\gamma}\kappa^{L_{\mathcal{R}}(\gamma^{\prime})}\prod_{\operatorname{cc}\in\chi}\prod_{f\in\mathcal{T}_{\operatorname{cc}}\sqcup\{e^{\prime}(\operatorname{cc})\}}(Q^{\ell^{\prime}(f)})_{x^{\prime}_{u}x^{\prime}_{v}}.$
For a fixed vector $\vec{\ell}=(\ell_{11},..,\ell_{it},...,\ell_{2m,T_{2m}})$ we have
@@
$\sum_{\begin{subarray}{c}\gamma^{\prime}\sim\gamma\\ \vec{\ell}(\gamma^{\prime})=\vec{\ell}\end{subarray}}q_{\mathrm{s}}(\gamma^{\prime})$ $\leq\kappa^{L_{\mathcal{R}}(\gamma^{\prime})}\sum_{\begin{subarray}{c}y_{u},x_{v}\in[N]\\ (u,v)\in\mathcal{Y}\times\mathcal{X}\end{subarray}}\prod_{f\in\mathcal{T}_{\operatorname{cc}}\sqcup\{e^{\prime}(\operatorname{cc})\}}(Q^{\ell(f)})_{x_{u}x_{v}}$
@@
$\leq N^{s-p}\delta^{L_{\mathcal{F}}}\kappa^{L_{\mathcal{R}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell(e(\operatorname{cc}))}{Q^{\ast}}^{\ell(e^{\prime}(\operatorname{cc}))}\right)$ (36)
where we applied ˜8.5. For a finite set $\mathcal{S}$ and an integer $I$ we introduce
$\mathcal{L}^{\prime}(\mathcal{S},I):=\left\{(\ell(s))_{s\in\mathcal{S}}\in[\ell]^{\mathcal{S}}:\penalty 10000\ \sum_{s\in\mathcal{S}}\ell(s)=I\right\}.$
We set $T:=\sum_{i=1}^{2m}T_{i}$ and $M=2m(\ell+1)-T$. We bound $|\mathcal{L}(L_{\mathcal{F}},\mathcal{F})|$ by $\ell^{F}$ and summing over all vector $\vec{\ell}$ possible we obtain
@@
$\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{s}}\leq N^{s-p}\ell^{F}\kappa\sum_{L_{\mathrm{d}}+L_{\mathcal{F}}+L_{\mathcal{R}}=M}(2\kappa)^{L_{\mathcal{R}}+R-1}\delta^{L_{\mathcal{F}}}\sum_{\begin{subarray}{c}L^{1}_{\mathrm{d}}+L^{2}_{\mathrm{d}}=L_{\mathrm{d}}\\ (\ell_{\operatorname{cc}}^{i})_{\operatorname{cc}\in\chi_{\mathrm{d}}}\in\mathcal{L}(L_{\mathrm{d}}^{i},\chi_{\mathrm{d}})\end{subarray}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}(Q^{\ell^{1}(\operatorname{cc})}{Q^{\ast}}^{\ell^{2}(\operatorname{cc})})$ (37)
where we used
$\sum_{(\ell(f))_{f\in\mathcal{R}}\in\mathcal{L}(L_{\mathcal{R}},\mathcal{R})}\kappa^{L_{\mathcal{R}}}=\binom{R-1}{L_{\mathcal{R}}+R-1}\kappa^{L_{\mathcal{R}}}\leq(2\kappa)^{L_{\mathcal{R}}+R-1}\kappa.$ (38)
For $0\leq L_{\mathrm{d}}^{1}+L_{\mathrm{d}}^{2}=L_{\mathrm{d}}$ fixed, denoting $\mathrm{x}_{\mathrm{d}}=|\chi_{\mathrm{d}}|$ we have
$\sum_{(\ell^{i}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{i})}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell^{1}_{\operatorname{cc}}}{Q^{\ast}}^{\ell^{2}_{\operatorname{cc}}}\right)$ $\leq\sum_{(\ell^{i}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{i}+\mathrm{x}_{\mathrm{d}})}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell^{1}_{\operatorname{cc}}}{Q^{\ast}}^{\ell^{2}_{\operatorname{cc}}}\right)$ (39)
$\leq\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{(\ell^{i}_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{i}+\mathrm{x}_{\mathrm{d}})}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{\ell^{1}_{t}}{Q^{\ast}}^{\ell_{t}^{2}}\right).$
Recall the definition of $\mathrm{b}_{\mathrm{x}}$ given by (9). We define the rest
$\mathrm{R}_{\mathrm{s}}:=\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{(\ell^{i}_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{i}+\mathrm{x}_{\mathrm{d}})}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(Q^{\ell^{1}_{t}}{Q^{\ast}}^{\ell_{t}^{2}}\right)-\Phi({\mathrm{b}}_{\mathrm{s}}^{L^{1}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}}{{\mathrm{b}}^{\ast}_{\mathrm{s}}}^{L^{2}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}})$
and we have
@@
$\mathrm{R}_{\mathrm{s}}$ $=\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{(\ell^{i}_{t})_{t}\in\mathcal{L}(T,L_{\mathrm{d}}^{i}+\mathrm{x}_{\mathrm{d}})}\sum_{k=1}^{T}\sum_{\begin{subarray}{c}V\subset T\\ \left|V\right|=k\end{subarray}}\frac{\delta^{\sum_{t\in V}\ell^{1}_{t}+\ell^{2}_{t}}}{N^{k}}\prod_{t\notin V}\operatorname{Tr}_{N}\left(Q_{\perp}^{\ell^{1}_{t}}{Q_{\perp}^{\ast}}^{\ell_{t}^{2}}\right)$
@@
$=\sum_{k=1}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{T=k}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{L^{i}_{\mathrm{d},k}\leq L^{i}_{\mathrm{d}}}\binom{T}{k}|\mathcal{L}^{\prime}(L^{1}_{\mathrm{d},k},k)||\mathcal{L}^{\prime}(L^{2}_{\mathrm{d},k},k)|\frac{\delta^{2\ell k}}{N^{k}}\sum_{\ast}\prod_{t=k+1}^{T}\operatorname{Tr}_{N}\left(Q_{\perp}^{\ell^{1}_{t}}{Q_{\perp}^{\ast}}^{\ell_{t}^{2}}\right)$
where the last summand is over all $(\ell_{t}^{i})_{t}\in\mathcal{L}(T-k,L^{i}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}-L_{\mathrm{d},k}^{i}-k)$ for $i\in\{1,2\}$. The last sum above verifies
$\left|\sum_{\ast}\prod_{t=k+1}^{T}\operatorname{Tr}_{N}\left(Q_{\perp}^{\ell^{1}_{t}}{Q_{\perp}^{\ast}}^{\ell_{t}^{2}}\right)\right|$ $\leq\sum_{\ast}\prod_{t=k+1}^{T}\left|\operatorname{Tr}_{N}\left(Q_{\perp}^{\ell^{1}_{t}}{Q_{\perp}^{\ast}}^{\ell_{t}^{1}}\right)\operatorname{Tr}_{N}\left(Q_{\perp}^{\ell^{2}_{t}}{Q_{\perp}^{\ast}}^{\ell_{t}^{2}}\right)\right|^{1/2}$
$\leq\sum_{\ast}\prod_{t=k+1}^{T}\|Q_{\perp}^{\ell^{1}_{t}}\|\|Q_{\perp}^{\ell^{2}_{t}}\|\leq(2\|Q_{\perp}\|)^{L_{\mathrm{d}}^{1}+L_{\mathrm{d}}^{2}+2\mathrm{x}_{\mathrm{d}}-L_{\mathrm{d},k}^{1}-L_{\mathrm{d},k}^{2}}$
where for the last two inequalities we applied Section˜6.4 and used that for $b,h\in\mathcal{A}$
$\Phi(bh^{*})\leq\sqrt{\tau(bb^{*})\tau(hh^{*})},\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \tau(b^{\ast}b)=\mathbb{E}[\|b^{\ast}b(e_{0})\|_{2}^{2}]\leq\|b\|^{2}.$
@@
The integer $n$ being fixed, for all integer $L=\mathrm{q}n+\mathrm{r}\geq 1$ with $0\leq\mathrm{r}\leq n-1$, We set $c(n):=(\kappa\vee 1)(1+\delta^{\prime})^{n}$ and we have
${\mathrm{N}}_{L,\mathrm{s}}^{L}=\|{\mathrm{b}}_{\mathrm{s}}^{L}\|\leq\|{\mathrm{b}}_{\mathrm{s}}^{n}\|^{\mathrm{q}}\|{\mathrm{b}}_{\mathrm{s}}\|^{\mathrm{r}}\leq{\mathrm{N}}_{n,\mathrm{s}}^{L-\mathrm{r}}(1+|\lambda_{2}(Q)|)^{n}\leq c(n){\mathrm{N}}_{n,\mathrm{s}}^{L}$ (40)
where we used that the operator norm is sub-multiplicative. Besides we have $|\mathcal{L}^{\prime}(L,k)|\leq\ell^{k}$ and therefore we obtain
$\mathrm{R}_{\mathrm{s}}$ $\leq 2m\ell c(n)^{2}\sum_{k=1}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{T=k}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\binom{T}{k}\left(\frac{\ell^{2}\delta^{2\ell}}{N}\right)^{k}\sum_{L_{\mathrm{d},k}\leq L_{\mathrm{d}}}{{\mathrm{N}}_{n,\mathrm{s}}}^{L_{\mathrm{d}}+2\mathrm{x}_{\mathrm{d},\mathrm{s}}-L_{\mathrm{d},k}}$
$\leq c(2m\ell)^{5}c(n)^{2}\frac{\rho_{n,\mathrm{s}}^{L_{\mathrm{d}}}}{N}.$
We used the condition (34) for the second inequality. Implementing the previous inequality in (39) we obtain
$\sum_{(\ell^{i}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{i})}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(Q^{\ell^{1}_{\operatorname{cc}}}{Q^{\ast}}^{\ell^{2}_{\operatorname{cc}}}\right)$ $\leq|\Phi({\mathrm{b}}_{\mathrm{s}}^{L^{1}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}}{{\mathrm{b}}^{\ast}_{\mathrm{s}}}^{L^{2}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}})|+\left|\mathrm{R}_{\mathrm{s}}\right|$
$\leq c(n)(2m\ell)^{5}\rho_{n,\mathrm{s}}^{L_{\mathrm{d}}}$ (41)
where we proceed as we did in (40) and obtained $|\Phi({\mathrm{b}}_{\mathrm{s}}^{L^{1}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}}{{\mathrm{b}}^{\ast}_{\mathrm{s}}}^{L^{2}_{\mathrm{d}}+\mathrm{x}_{\mathrm{d}}})|\leq c(n)(\mathrm{N}_{n,\mathrm{s}}\vee 1)^{L_{\mathrm{d}}}$. For the convex combination case we introduce
$\mathrm{S}_{\mathrm{c}}:=(1-r)^{2(\mathrm{x}_{\mathrm{d}}-1)}\sum_{(\ell_{\operatorname{cc}},\ell^{\prime}_{\operatorname{cc}})_{\operatorname{cc}}}\prod_{\operatorname{cc}\in\chi_{\mathrm{d}}}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{\operatorname{cc}}}\{rQ^{\ast}\}^{\ell_{\operatorname{cc}}^{\prime}}\right)$
and we have
$\mathrm{S}_{\mathrm{c}}$ $\leq\sum_{T=0}^{L_{\mathrm{d}}^{1}\land L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}}\sum_{(\ell^{i}_{\operatorname{cc}})_{\operatorname{cc}}\in\mathcal{L}^{\prime}(\chi_{\mathrm{d}},L_{\mathrm{d}}^{i})}(1-r)^{2(T-1)}\prod_{t=1}^{T}\operatorname{Tr}_{N}\left(\{rQ\}^{\ell_{t}^{1}}\{rQ^{\ast}\}^{\ell_{t}^{2}}\right)$
$\leq|\Phi({\mathrm{b}}_{\mathrm{c}}^{L_{\mathrm{d}}^{1}+\mathrm{x}_{\mathrm{d}}}{{\mathrm{b}}_{\mathrm{c}}^{\ast}}^{L_{\mathrm{d}}^{2}+\mathrm{x}_{\mathrm{d}}})|+|\mathrm{R}_{\mathrm{c}}|\leq c(n)(2m\ell)^{5}\rho_{n,\mathrm{c}}^{L_{\mathrm{d}}}$
where $\mathrm{R}_{\mathrm{c}}$ is defined analogously than for $\mathrm{x}=\mathrm{s}$. Back to the case $\mathrm{x}=\mathrm{s}$, implementing the inequality (41) into (37) we obtain
$\sum_{\gamma^{\prime}\sim\gamma}q_{\mathrm{s}}(\gamma^{\prime})$ $\leq N^{s-p}\ell^{F}(2m\ell)^{6}c(n)^{2}\sum_{L_{\mathcal{R}}+L_{\mathcal{F}}+L_{d}=2m\ell}(2\kappa)^{L_{\mathcal{R}}}\delta^{L_{\mathcal{F}}}\rho_{n,\mathrm{s}}^{L_{\mathrm{d}}}$
$\leq N^{s-p}\ell^{F}(2m\ell)^{8}c(n)^{2}\rho_{n,\mathrm{s}}^{2m\ell}$
@@
where we used the fact that owing to ˜6.4 we have $(2\kappa)\vee\delta\vee\mathrm{N}_{n,\mathrm{s}}\geq 1$ and the fact that $M+R+\mathrm{x}_{\mathrm{d}}\leq 2m(\ell+1)$. It remains to apply ˜8.6 and adapt the constant $c(n)>0$.
We now give the proof of ˜8.6.
Proof of ˜8.6.
@@
[…]